OpenStudy (anonymous):

The difference between any two consecutive interior angles of a polygon is 5°. If the smallest angle is 120°, find the number of the sides of the polygon. solving this, values of 'n' come as either 9 or 16. what to do ? thanks

OpenStudy (jiteshmeghwal9):

is 'n' the number of vertices??

OpenStudy (anonymous):

n= number of sides..which is equal to vertices as well

OpenStudy (anonymous):

sum of internal angles of polygon = 180 x (n-2)

OpenStudy (anonymous):

it can b plugged for n=16 as well, the answer is actually 9 , but why?

OpenStudy (anonymous):

one said that the naswer can be both..may be polygon can be formed in both the cases?

OpenStudy (jiteshmeghwal9):

first of all find the angles of a polygon let the bigger angle be 'x'. then A/q x-120=5 x=5+120=125

OpenStudy (anonymous):

if you use the equation given above byjitesh then you should be getting 9

OpenStudy (anonymous):

jitesh what is A/q?